Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 421
Entry time: 24 May 2023, 09:50
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 24.05.2023 |
|
|
General Announcement:
-
CTRL:
- Not present.
DA:
- Nothing to report.
EEE:
- Nothing to report.
VAC:
- Nothing to report.
XPD:
- Everything working fine.
DET:
- Not present.
ITDM:
- Nothing to report.
XRO:
- Nothing to report.
LAS:
- PP-lasers are back and running. All fine.
UNSYS:
- Nothing to report.
SPB:
- Nothing to report.
FXE:
- Nothing to report.
MID:
- Issue with clipping the beam was found and relates to a combination of devices.
- Progress is being made for user experiment.
HED:
- Preparation for next week ongoing.
- Needs to get in contact with RCs for next Monday tuning.
SCS:
- Preparing the THz setup for beamtime.
SQS:
- Nothing to report.
SXP:
- Used beam yesterday for commissioning KB. |